# HER2 and FISH Status Prediction in Breast Biopsy H&E-Stained Images Using Deep Learning

### v2

@@ Abstract
The current standard for detecting human epidermal growth factor receptor 2 (HER2) status in breast cancer patients relies on HER2 amplification, identified through fluorescence in situ hybridization (FISH) or immunohistochemistry (IHC). However, hematoxylin and eosin (H&E) tumor stains are more widely available, and accurately predicting HER2 status using H&E could reduce costs and expedite treatment selection. Deep Learning algorithms for H&E have shown effectiveness in predicting various cancer features and clinical outcomes, including moderate success in HER2 status prediction. In this work, we employed a customized weak supervision classification technique combined with MoCo-v2 contrastive learning to predict HER2 status. We trained our pipeline on 182 publicly available H&E Whole Slide Images (WSIs) from The Cancer Genome Atlas (TCGA), for which annotations by the pathology team at Yale School of Medicine are publicly available. Our pipeline achieved an Area Under the Curve (AUC) of 0.85 ± 0.02 across four different test folds. Additionally, we tested our model on 44 H&E slides from the TCGA-BRCA dataset, which had an HER2 score of 2+ and included corresponding HER2 status and FISH test results. These cases are considered equivocal for IHC, requiring an expensive FISH test on their IHC slides for disambiguation. Our pipeline demonstrated an AUC of 0.81 on these challenging H&E slides. Reducing the need for FISH test can have significant implications in cancer treatment equity for underserved populations.
@@ I Introduction
TABLE I: Scoring guidelines for Human Epidermal Growth Factor Receptor 2 (HER2) by the American Society of Clinical Oncology/College of American Pathologists. [8]
@@ II Related Work
Several studies have explored the use of computational methods and deep learning algorithms to enhance the detection and classification of HER2 status in breast cancer, particularly using hematoxylin and eosin (H&E) stained slides and immunohistochemistry (IHC). One study [9] developed a pipeline combining stain separation and deep neural networks to classify HER2 status from H&E-stained whole slide images, demonstrating the potential for automated quantification of mutation-specific protein overexpression in digital pathology. Another approach utilized a convolutional neural network (CNN) trained on manually annotated H&E slides, achieving high accuracy in predicting HER2 status and treatment response, suggesting that H&E-based algorithms can effectively support clinical evaluations [15]. In one more study, a GrayMap+ CNN model was employed to predict HER2 expression levels and gene status directly from IHC slides, showing strong agreement with pathologist assessments and high accuracy in classification [18]. In another approach, deep learning algorithms have been applied to HER2 2+ IHC slides to predict gene amplification status, achieving moderate accuracy but highlighting the potential for triaging cases in settings where in situ hybridization testing is not readily available [19]. Another work developed a 3-block-DenseNet deep learning model to predict HER2 expression from ultrasound images providing a non-invasive method for HER2 prediction [20].
III Methodology
@@
The task at hand involves classifying whole slide images stained by inexpensive and ubiquitous hematoxylin and eosin (H&E) stains. Because we only know the overall HER2 status of a patient but cannot localize the tissue regions showing HER2 positivity in H&E slides, we opted for using a weakly supervised approach. Weakly supervised classification tasks in pathology typically involve training sets with known labels for each Whole Slide Image (WSI) but lack class-specific annotations or information for individual pixels or regions. Attention-based Multiple Instance Learning (MIL) [10] with clustering extends the MIL [11] framework, which is well-suited for multi-class classification. Traditional MIL utilizes a non-trainable aggregation function, such as max pooling, where the slide-level prediction is determined by the patch exhibiting the highest prediction probability. In contrast, attention-based MIL with clustering employs a trainable and interpretable attention-based pooling function. This function aggregates slide-level representations from individual patch-level representations. Within the attention-based pooling mechanism, the attention network generates two distinct sets of attention scores, each corresponding to the classes in the binary classification problem. In our study, we classified breast cancer whole slide images into HER2 positive and HER2 negative categories using a weakly supervised technique proposed by Fremond et al [12].
@@
The deep learning model proposed by Fremond et al [12] encompasses a comprehensive approach that begins with pre-processing the whole-slide images to prepare them for detailed analysis. It then trains a feature extractor model using contrastive self-supervised learning, specifically MoCo-v2 [13], aimed at learning patch-level morphological features. Subsequently, all patch-level feature vectors are extracted from the self-supervised learning encoder. The process culminates in the training and inference of an attention-based classification model [14], which molecularly classifies whole-slide images using the gathered patch-level feature vectors.
@@
During the pre-processing phase, the tissue in each whole-slide image was automatically segmented using Otsu thresholding. The identified tissue regions were then divided into non-overlapping square patches, each measuring $360\,\mu m$ at 40×40$\times$40 × magnification, and resized to $224\times 224$ pixels. For the training set, which consisted of $120$ whole-slide images, there was an average of $1800$ patches per image. To ensure quality, patches with minimal or no tissue were excluded by applying a minimum threshold of $20$ to the median value of each $8$-bit RGB channel. Additionally, these patches were filtered out if their mean pixel intensity across the RGB channels was approximately 240 ± 15 using a specific technique [25]. A Whole Slide Image and its corresponding patch grid, derived through the preprocessing method, are depicted in Figure 2.
@@
To train the feature extractor model using self-supervised learning, we curated a dataset of image patches by sampling from each whole-slide image. To ensure fair representation of each molecular class, we strategically sampled an optimal number of patches from whole-slide images corresponding to the two classes involved in our study. This methodology produced a dataset consisting of 216,000 patches derived from 120 whole-slide images, with an even distribution across the two classes. We then trained MoCo-v2 [13] for 300 epochs using a ResNet-50 encoder on these Whole Slide Image patches. MoCo-v2 is a contrastive learning framework designed to discern between similar and dissimilar data pairs, which it approaches as a dictionary lookup challenge. The learning rate was set at 0.06 with a temperature parameter of 0.07. Central to this framework is the InfoNCE loss function [13](equation represented below), which effectively measures the contrast between data pairs. In this setup, q denotes a query image, $x^{+}$ represents the positive (similar) key sample, and $x^{-}$ includes the negative (dissimilar) key samples. The temperature hyperparameter, $\tau$ , adjusts the sensitivity of the loss function. Within the framework’s instance discrimination pretext task, a positive pair consists of a query and a key that are data-augmented versions of the same image, while a negative pair is formed when the query and key derive from different images. This methodology aids in developing features that are resilient to variations introduced by data augmentation, enhancing the model’s generalization capabilities. The data augmentation techniques applied during self-supervision training included random rotations, vertical and horizontal flips, color jitter, and Gaussian Blur.
@@
Feature vectors of size 2048 were obtained from each patch of the whole-slide images using a self-supervised learning ResNet-50 encoder at its final layer. Subsequently, these vectors for all patches within a single whole-slide image were aggregated. Following this, an attention-based slide classifier, derived from the CLAM [14] architecture but without the secondary clustering objective, was trained over 100 epochs using cross entropy loss. The whole training pipeline is depicted in Figure  3.

### v3

@@ Abstract
The current standard for detecting human epidermal growth factor receptor 2 (HER2) status in breast cancer patients relies on HER2 expression identified through immunohistochemistry (IHC) or amplification identified through fluorescence in situ hybridization (FISH). However, hematoxylin and eosin (H&E) tumor stains are more widely available, and accurately predicting HER2 status using H&E could reduce costs and expedite treatment selection. Deep Learning algorithms for H&E have shown effectiveness in predicting various cancer features and clinical outcomes, including moderate success in HER2 status prediction. In this work, we employed a customized weak supervision classification technique combined with MoCo-v2 contrastive learning for self-supervised feature extraction training to predict HER2 status. We trained our pipeline on 182 publicly available H&E whole slide images (WSIs) from The Cancer Genome Atlas (TCGA), for which annotations by the pathology team at Yale School of Medicine are publicly available. Our pipeline achieved an Area Under the Curve (AUC) of 0.85 ± 0.02 across four different test folds. Additionally, we tested our model on 44 H&E slides from the TCGA-BRCA dataset, which had an HER2 score of 2+ and included corresponding HER2 status and FISH test results. These cases are considered equivocal for IHC, requiring an expensive FISH test on their IHC slides for disambiguation. Our pipeline demonstrated an AUC of 0.81 on these challenging H&E slides. Reducing the need for FISH test can have significant implications in cancer treatment equity for underserved populations.
@@ I Introduction
TABLE I: Scoring guidelines for human epidermal growth factor receptor 2 (HER2) by the American Society of Clinical Oncology/College of American Pathologists (ASCO/CAP). [8]
@@ II Related Work
Several studies have explored the use of computational methods and deep learning algorithms to enhance the detection and classification of HER2 status in breast cancer, particularly using hematoxylin and eosin (H&E) stained slides and immunohistochemistry (IHC). One significant approach was presented by Anand et al.  [9], who developed a multi-stage deep learning pipeline to estimate HER2 status from H&E stained images. Their method combined a stain separation module with three convolutional neural networks (CNNs) to classify HER2-positive and HER2- negative samples. Another approach utilized a convolutional neural network (CNN) trained on manually annotated H&E slides, achieving high accuracy in predicting HER2 status and treatment response, suggesting that H&E-based algorithms can effectively support clinical evaluations [15]. In one more study, a GrayMap+ CNN model was employed to predict HER2 expression levels and gene status directly from IHC slides, showing strong agreement with pathologist assessments and high accuracy in classification [18]. In another approach, deep learning algorithms have been applied to HER2 2+ IHC slides to predict gene amplification status, achieving moderate accuracy but highlighting the potential for triaging cases in settings where in situ hybridization testing is not readily available [19]. Another work developed a 3-block-DenseNet deep learning model to predict HER2 expression from ultrasound images providing a non-invasive method for HER2 prediction [20].
@@ III Methodology
The task at hand involves classifying whole slide images stained by inexpensive and ubiquitous hematoxylin and eosin (H&E) stains. Because we only know the overall HER2 status of a patient but cannot localize the tissue regions showing HER2 positivity in H&E slides, we opted for using a weakly supervised approach. Weakly supervised classification tasks in pathology typically involve training sets with known labels for each whole slide image (WSI) but lack class-specific annotations or information for individual pixels or regions. Attention-based Multiple Instance Learning (MIL) [10] with clustering extends the MIL [11] framework, which is well-suited for multi-class classification. Traditional MIL utilizes a non-trainable aggregation function, such as max pooling, where the slide-level prediction is determined by the patch exhibiting the highest prediction probability. In contrast, attention-based MIL with clustering employs a trainable and interpretable attention-based pooling function. This function aggregates slide-level representations from individual patch-level representations. Within the attention-based pooling mechanism, the attention network generates two distinct sets of attention scores, each corresponding to the classes in the binary classification problem. In our study, we classified breast cancer whole slide images into HER2 positive and HER2 negative categories using a weakly supervised technique proposed by Fremond et al. [12].
@@
The deep learning model proposed by Fremond et al. [12] encompasses a comprehensive approach that begins with pre-processing the whole slide images to prepare them for detailed analysis. It then trains a feature extractor model using contrastive self-supervised learning, specifically MoCo-v2 [13], aimed at learning patch-level morphological features. Subsequently, all patch-level feature vectors are extracted from the self-supervised learning encoder. The process culminates in the training and inference of an attention-based classification model [14], which molecularly classifies whole slide images using the gathered patch-level feature vectors.
@@
During the pre-processing phase, the tissue in each whole slide image was automatically segmented using Otsu thresholding. The identified tissue regions were then divided into non-overlapping square patches, each measuring $360\,\mu m$ at 40×40$\times$40 × magnification, and resized to $224\times 224$ pixels. To ensure quality, patches with minimal or no tissue were excluded by applying a minimum threshold of $20$ to the median value of each $8$-bit RGB channel. Additionally, patches with a mean pixel intensity across the RGB channels of approximately 245 ± 10, indicating near-white pixels, were filtered out using a specific technique[25]. A whole slide image and its corresponding patch grid, derived through the preprocessing method, are depicted in Figure 2.
@@
To train the feature extractor model using self-supervised learning, we curated a dataset of image patches by sampling from each whole slide image. To ensure fair representation of each molecular class, we strategically sampled an optimal number of patches from whole slide images corresponding to the two classes involved in our study. This approach generated a dataset of $240,000$ tissue region image patches extracted from 120 whole slide images, evenly distributed between the two classes. We then trained MoCo-v2 [13] for 300 epochs using a ResNet-50 encoder on these whole slide image patches. MoCo-v2 is a contrastive learning framework designed to discern between similar and dissimilar data pairs, which it approaches as a dictionary lookup challenge. The learning rate was set at 0.06 with a temperature parameter of 0.07. Central to this framework is the InfoNCE loss function [13](equation represented below), which effectively measures the contrast between data pairs. In this setup, q denotes a query image, $x^{+}$ represents the positive (similar) key sample, and $x^{-}$ includes the negative (dissimilar) key samples. The temperature hyperparameter, $\tau$ , adjusts the sensitivity of the loss function. Within the framework’s instance discrimination pretext task, a positive pair consists of a query and a key that are data-augmented versions of the same image, while a negative pair is formed when the query and key derive from different images. This methodology aids in developing features that are resilient to variations introduced by data augmentation, enhancing the model’s generalization capabilities. The data augmentation techniques applied during self-supervision training included random rotations, vertical and horizontal flips, color jitter, and Gaussian Blur.
@@
Feature vectors of size 2048 were obtained from each patch of the whole slide images using the self-supervised ResNet-50 encoder at its final layer. Subsequently, these vectors for all patches within a single whole slide image were aggregated. Following this, an attention-based slide classifier, derived from the CLAM [14] architecture but without the secondary clustering objective, was trained over 100 epochs using cross entropy loss. The whole training pipeline is depicted in Figure  3.